To sterilize a 50.0-g glass babby bottle, we must raise its temperature from 22.0 degree celcius to 95.0 degree celcius . How much heat transfer is required?


1
Expert's answer
2021-11-29T11:41:58-0500

"Q=cm(t_2-t_1)=600\cdot 0.05\cdot(95-22)=2190\ (J)" . Answer